OpenStudy (anonymous):

if a,b,c are positive and a^-2b^2c^2 > a^-3b^3c^2 which of the following must be true? 1. a>b 2. a>c 3. b>c I'd really appreciate the help :')

OpenStudy (callisto):

\[a^{-2}b^2c^2 > a^{-3}b^3c^2 \]\[\frac{b^2c^2}{a^2} > \frac{b^3c^2}{a^3} \]Multiply a^3 to both sides, then divide both sides by b^2c^2 to get your answer.

OpenStudy (anonymous):

So you get b^3 c^2?

OpenStudy (anonymous):

How do I know which one is greater ? 1,2, or 3 though?

OpenStudy (callisto):

I'm sorry that I was away. When you multiply a^3 to both sides of the question, what do you get?

OpenStudy (anonymous):

a^3b^2C^2/a^2 > b^3c^2

OpenStudy (anonymous):

its fine!

OpenStudy (callisto):

For the left side, can you simplify the expression?

OpenStudy (anonymous):

Yeah and then you get a^3b^2c^2/a^2 = 1/b^2c^2 and then you can cancel the equation out

OpenStudy (callisto):

That doesn't look right to me ... \[a^{-2}b^2c^2 > a^{-3}b^3c^2\]\[\frac{b^2c^2}{a^2} > \frac{b^3c^2}{a^3}\]Multiply both sides by a^3\[a^3 \times \frac{b^2c^2}{a^2} > a^3 \times\frac{b^3c^2}{a^3}\]Cancel the common factor on both sides, what do you get?

OpenStudy (anonymous):

Omg can you just tell me the answer I hate math! but you a^3 X b^2c^2/a^2 = b^3C^2

OpenStudy (callisto):

I can't, I'm sorry! \[a^3 \times \frac{b^2c^2}{a^2} > \cancel{a^3} \times\frac{b^3c^2}{\cancel{a^3}}\]\[a^3 \times \frac{b^2c^2}{a^2} > b^3c^2\]That's what you got. For the left, you can do something similar. \[a(a^2) \times \frac{b^2c^2}{a^2} > b^3c^2\] Can you try again? Cancel the common factor on the left.

OpenStudy (anonymous):

This would be easier if I could type like that but okay a^3b^2c^2 >b^3c^2

OpenStudy (callisto):

You can use the equation button to type that :D Still... it's not right :| \[a(a^2) \times \frac{b^2c^2}{a^2} > \cancel{a^3} \times\frac{b^3c^2}{\cancel{a^3}}\]\[a\cancel{(a^2)} \times \frac{b^2c^2}{\cancel{a^2}} > \cancel{a^3} \times\frac{b^3c^2}{\cancel{a^3}}\]What's left?

OpenStudy (anonymous):

Omg lol im so dumb \[ab^2c^2 > b^3c^2 ? \]

OpenStudy (callisto):

It looks right now :) Now, divide both side by c^2. What do you get?

OpenStudy (anonymous):

\[ab^2> b^3\]

OpenStudy (callisto):

Yup! Now, divide both sides by b^2, what do you get?

OpenStudy (anonymous):

a>b ??

OpenStudy (callisto):

Yup :)

OpenStudy (callisto):

You did it !! :)

OpenStudy (anonymous):

Omg thankyou so much!! :')

OpenStudy (callisto):

Welcome :D

OpenStudy (anonymous):

you're great aahh thanks <3333